OpenStudy (anonymous):

find an equation of the line that goes through the point (2,1) and has a slope of 4. leave equation in point-slope form

OpenStudy (anonymous):

You use the point-gradient formula for a straight line.\[y-y_1=m(x-x_1)\]Here,\[y-1=4(x-2)\]

OpenStudy (anonymous):

\[y=4x-1\]